# ISMRNN: An Implicitly Segmented RNN Method with Mamba for Long-Term Time Series Forecasting

### v4

@@ 1 Introduction
Long-term series forecasting is the process of predicting the future values of a sequence over an extended period. This task primarily aims to forecast long-term trends, which hold significant importance in domains such as climate prediction (Aghelpour et al., 2019), decision support (Nooruldeen et al., 2022), and policy formulation (Zhang et al., 2022). Long-term sequence data often exhibit higher levels of uncertainty, which can result in reduced prediction accuracy. Additionally, longer forecasting horizons require models to incorporate a broader historical context for accurate predictions, thereby increasing the complexity of modeling. Recent advancesments in long-term sequence analysis have shifted towards deep learning methods, highlighting the crucial need to develop diverse and effective time-series methodologies.
Long-term time series analysis and forecasting comprise two critical components: (1) trend identification and analysis, which involves recognizing long-term directional changes in the data, This component is crucial for understanding the behavior of time series, enabling the comprehension of past patterns and the prediction of potential future developments. (2) Uncertainty and noise management: Due to the extended time range of long-term time series data, they often exhibit increased levels of uncertainty and noise. Effectively managing and mitigating these uncertainties and noises is vital for improving the accuracy of predictions.
@@
In the context of long-term time series analysis, it is essential to capture the sequential relationships between data. Commonly adopted feature representation methods in time series analysis include Convolutional Neural Networks (Wang et al., 2022; Wu et al., 2022; Foumani et al., 2021), Multilayer Perceptrons (Das et al., 2023a; Chen et al., 2023; Zeng et al., 2023), attention-based encoding layers (Das et al., 2023b; Nie et al., 2022; Zhang and Yan, 2022), and Recurrent Neural Networks networks based on sequence structure (Lin et al., 2023). CNN-based methods use convolutional kernels to capture temporal and spatial sequences (Foumani et al., 2021); however, the limited size of these kernels prevent the model from effectively capturing long-range dependencies in time series, thus limiting the expressiveness of the feature representation. Transformer-based methods, leveraging attention mechanisms to improve the capture of spatiotemporal sequences, thereby enhancing feature representation capabilities. MLPs have been exrensively applied in time series prediction, with models such as DLinear (Zeng et al., 2023), which use simple linear layers and channel-independent methods, achieving superior performance compared to the advanced Transformer architectures of its time, prompting a reevaluation of the need for complex Transformer structures in time series prediction.
The Mamba architecture, proposed by Gu and Dao(Gu and Dao, 2023), constitutes a state space model tailored for efficient sequence modeling. It integrates parameterized matrices and a hardware-aware parallel computing algorithm, achieving impressive results in tasks such as language modeling and time series forecasting. Mamba is notable for its ability to capture long-term dependencies while maintaining linear computational complexity, making it a strong competitor to Transformer-based models. This capability is facilitated through its adept management of state space models and the implementation of a selective state space mechanism designed to optimize performance.
@@
RNN-based methods have achieved remarkable success in long-term time series analysis tasks. The inherent sequential architecture of RNNs enables them to handle short-term time series prediction tasks effectively. However, as the length of the predicted sequence increases, the model’s prediction error rapidly escalates with the propagation of features through the RNN recurrent units (Yunpeng et al., 2017). To address this limitation, Shengsheng Lin et al. (Lin et al., 2023) proposed the segmentation-based SegRNN model. This model aims to minimize the number of RNN recurrent iterations while retaining the sequential information, achieving significant success in long-term time series foreasting tasks. SegRNN employs a segmented approach to time series, which may disrupt the continuity of the sequences. Although the segmentation strategy effectively balances sequence information and iteration count, information loss within the recurrent structure occurs as the prediction horizon extends, thereby limiting the model’s capability.
In this work, our objective is to further reduce information loss within the recurrent structure while introducing implicit segmentation to enhance the segmentation strategy. Specifically, we propose a long-term time series forecasting model that utilizes the Mamba structure, implicit segmentation, and residual blocks, termed ISMRNN, the main contributions of ISMRNN can be summarized as follows:
Introducing the Mamba structure to effectively preprocessing the information.
@@
The implementation of residual network in the encoding phase reduces the accumulation of errors during encoding stage.
@@ 2.1 Transformer Models
The core of applying Transformer models to long-term time series tasks lies in their ability to capture long-term dependencies using the attention mechanism (Hao and Cao, 2020). The general form of the self-attention mechanism is given by:
Significant research efforts have advanced the application of Transformer architectures in long-term time series analysis. For instance, PatchTST(Nie et al., 2022) introduces a method of segmenting time series into patches, reducing computational complexity and improving the model’s ability to capture local semantic information. It also employs self-supervised learning, masking random patches and training the model to reconstruct them, enhancing forecasting accuracy. iTransformer(Liu et al., 2023) addresses scalability by integrating hierarchical attention mechanisms and temporal convolutional networks, capturing both short-term and long-term dependencies more effectively. This lightweight architecture balances accuracy and efficiency, making it suitable for real-world applications. Crossformer (Zhang and Yan, 2022) employs a Dimension-segment-wise structure to embed time series into a two-dimensional vector, effectively capturing dependencies across time and dimensions through its Two-stage-attention structure. The segmentation technique utilized in Crossformer is particularly advantageous for long-term time series forecasting.
@@ 2.2 MLP Models
TiDE (Das et al., 2023b) introduces a residual structure within the linear layer to encode past time series and covariates, subsequently decoding the encoded time series with future covariates to obtain prediction results. The core operation of TiDE involves projecting features and encoding them through residual blocks, which can be described as:
where and are the weights and biases of the projection layer, respectively. The ResBlock operation encapsulates the residual connections used to maintain the integrity of the encoded features while integrating the covariates.
MLP-based methods for long-term time series analysis have demonstrated outstanding predictive accuracy while maintaining low time-space complexity, thus significantly advancing the development of MLP in long-term time series forecasting.
@@ 2.3 CNN Models
The core of time series analysis based on CNN architectures lies in utilizing convolutional kernels to extract temporal and channel features. MICN (Wang et al., 2022) aims to fully exploit the latent information in time series by designing a multi-scale hybrid decomposition module to decompose the complex patterns of input time series. Each module within this design leverages subsampling convolution and equidistant convolution to respectively extract local characteristics and global correlations. TimesNet (Wu et al., 2022) transforms one-dimensional time series into a set of two-dimensional tensors based on multiple periodicities. It further introduces the TimesBlock to adaptively uncover these multi-periodic patterns and effectively extracts intricate temporal dynamics from the transformed tensors through parameter-efficient initializations. These methodologies have attained remarkable results in the domain of long-term time series forecasting.
2.4 SegRNN Model
As the only RNN-based model specifically designed for long-term time series tasks, SegRNN has achieved remarkable performance. SegRNN addresses the challenge of retaining sequence information while minimizing information loss within the recurrent structure through two key methods:
SegRNN replaces point-wise iterations with segment-wise iterations. This approach balances the trade-off between sequence information retention and the number of recurrent iterations, leading to improved outcomes.
SegRNN replaces recurrent multi-step forecasting with parallel multi-step forecasting. This parallel decoding approach enhances the model’s predictive capability and inference speed.
By combining these two strategies, SegRNN achieves excellent performance within a lightweight model structure, along with superior inference speed.
@@ 3.1 Overview
Previous work, such as SegRNN, has achieved remarkable results by employing segmentation and parallel decoding techniques. However, the fixed segmentation approach restricts the model’s capacity and disrupts data continuity. Additionally, segmentation methods still face the challenge of cumulative information processing within the recurrent structure.
@@
3.2 Mamba structure
The Mamba architecture, initially proposed by Gu and Dao (2023), has demonstrated outstanding performance in various fields such as large language models and long time series forecasting. By setting model parameters to be content-related, the Mamba architecture achieves powerful data selection capabilities, which are crucial for time series forecasting tasks. Therefore, we utilize the Mamba architecture at the initial stage to preprocess and enhance our time series data. It is important to note that our model does not incorporate the convolutional layer of the Mamba architecture except for weather dataset. This difference is mainly based on performance considerations, as we believe that local information can be effectively captured through segmentation.
@@ 3.3 Implicit Segmentation
The segmentation method employed by SegRNN, which we refer to as explicit segmentation to distinguish it from our segmentation approach, is implemented by truncating the time dimension. For a given time series , selecting a segmentation length results in segmented vectors of size , where represents the number of segments.
@@
To address these issues, we propose an implicit segmentation strategy, as illustrated in Figure 2. This strategy initially decomposes the time series through linear mapping and subsequently transforms the information into the embedded hidden state . Notably, after segmentation, each segment includes both the information of that segment and some additional information, ensuring more continuous information processing and thereby enhancing the model’s performance.
@@ 3.4 Residual Structure
SegRNN aims to minimize the number of iterations in the RNN while preserving sequential information. Extensive experiments show that increasing the number of iterations negatively impacts information retention. SegRNN uses data segmentation to reduce the data’s flow through the RNN structure. Specifically, for the vector obtained through segmentation in SegRNN, , we transform it in the RNN structure with a hidden layer dimension using the following formula:
@@
Using segmentation techniques, SegRNN reduces the number of iterations from to . However, the number of iterations still increases with the forecasting time steps, which limits the model’s capability due to the loss of information during these iterations. To address this issue, we incorporate a residual structure into the model. For the segmented vector , we apply a linear mapping to , resulting in . Finally, we add to the corresponding position in to obtain the final output vector from the encoder.
@@ 4 Experiments
Our experiments are structured as follows. We first introduce the datasets used for long-term time series analysis. Subsequently, we provide a detailed description of the experimental setup and baseline models for comparison. Then we presents and analyzes the performance metrics. We further conducted ablation studies to investigate the effectiveness of the model. Finally, we investigated the efficiency of the model, particularly focusing on its memory consumption to demonstrate its lightweight characteristic. All experiments in this section were conducted on two NVIDIA T4 GPU, The source code is available at https://github.com/UniqueoneZ/ISMRNN.git.
@@ 4.1 Datasets
The evaluation is conducted across seven real-world datasets spanning various domains, including power transformer temperature, weather, electricity load, with channel ranging from 7 to 321 and frequencies from every 10 minutes to hourly recordings. Detailed dataset information is provided in Table 1.
@@ 4.3 Main Result
We selected two evaluation metrics, MSE and MAE, with prediction steps of 96, 192, 336, and 720 time steps. The resulting forecasts are presented in Table 2. Note that all data, except for our method, originate from official code repositories and original papers.
@@ 4.4 Ablation study
We further conducted an ablation study to verify the effectiveness of incorporating the Mamba structure, implicit segmentation, and residual structures into our model. We performed experiments excluding the Mamba structure and implicit segmentation structure respectivly. For clarity, in the table below, "M" denotes preprocessing using the Mamba structure, and "LR" represents the addition of linear segmentation and residual structure. The results are shown in Table 3:
@@ 4.5 Model efficiency
As shown in the figure, compared to SegRNN, our model only increases memory usage and training time. However, in contrast to some transformer-based models, our method still holds significant spatiotemporal advantages.
@@ 5 Conclusion
In this work, we present the ISMRNN method: a network that incorporates Mamba structure, implicit segmentation and residual structure into the SegRNN framework. These enhancements mitigate the information loss caused by fixed segmentation in time series while enhance the information from the recurrent structure of the model. It is noteworthy that our proposed implicit segmentation method is not limited to a specific long-term time series forecasting method.
We further conducted experiments on six real-world benchmark datasets for long time series. The results demonstrate that our model yields significant imporvements in some real-world scenarios. To validate the efficiency of our method in terms of space complexity, we compared the memory consumption of ISMRNN against SegRNN on datasets of varying scales. The results suggest that our method maintains high space efficiency, providing a solid foundation for the model’s application.
Future work will explore how these structures enhance the effect of the model and identify the particular characteristics of time series where the ISMRNN method excels with corresponding empirical research.

### v5

@@ 1 Introduction
Long-term series forecasting is the process of predicting the future values of a sequence over an extended period. This task primarily aims to forecast long-term trends, which hold significant importance in domains such as climate prediction (Aghelpour et al., 2019), decision support (Nooruldeen et al., 2022), and policy formulation (Zhang et al., 2022). Long-term sequence data often exhibit higher levels of uncertainty, which results in reduced prediction accuracy. Additionally, longer forecasting horizons require models to incorporate a broader historical context for accurate predictions, thereby increasing the complexity of modeling. Recent advancesments in long-term sequence analysis have shifted towards deep learning methods, highlighting the crucial need to develop diverse and effective time-series methodologies.
@@
Long-term time series analysis involves identifying trends to understand past and predict future changes, managing uncertainty and noise to enhance prediction accuracy, and capturing the sequential relationships within the data. Commonly adopted feature representation methods in time series analysis include Convolutional Neural Networks(CNNs) (Wang et al., 2022; Wu et al., 2023; Foumani et al., 2021), Multilayer Perceptrons(MLPs) (Chen et al., 2023; Zeng et al., 2023; Das et al., 2023; Li et al., 2023), attention-based Transformer models (Nie et al., 2023; Zhang and Yan, 2022; Zhou et al., 2021, 2022; Wu et al., 2021; Liu et al., 2021), and Recurrent Neural Networks(RNNs) based on sequence structure (Lin et al., 2023; Salinas et al., 2020). CNN-based methods use convolutional kernels to capture temporal and spatial sequences (Foumani et al., 2021); however, the limited size of these kernels prevent the model from effectively capturing long-range dependencies in time series, thus limiting the expressiveness of the feature representation. Transformer-based methods, leveraging attention mechanisms to improve the capture of spatiotemporal sequences, thereby enhancing feature representation capabilities. MLPs have been exrensively applied in time series prediction, with models such as DLinear (Zeng et al., 2023), which use simple linear layers and channel-independent methods, achieving superior performance compared to the advanced Transformer architectures of its time, prompting a reevaluation of the need for complex Transformer structures in time series prediction.
@@
RNN-based methods have achieved remarkable success in long-term time series analysis tasks. The sequential structure of RNNs allows them to effectively handle short-term time series prediction tasks. However, as the length of the prediction sequence increases, traditional RNN structures gradually become ineffective due to their inability to handle long-term dependencies and gradient issues (Yunpeng et al., 2017). To address this limitation, Shengsheng Lin et al. (Lin et al., 2023) proposed the SegRNN model. By segmenting the time series and employing a parallel decoding strategy, SegRNN minimizes the number of RNN recurrent iterations while retaining sequential information to the greatest extent. This approach has led to significant success in both the accuracy and computational efficiency of long-term time series forecasting tasks. Nevertheless, SegRNN has several limitations: The segmentation process disrupts the continuity of the sequences, and information across different segments is not fully utilized. Additionally, the segmentation strategy effectively balances sequence information with iteration count; however, information loss within the recurrent structure increases as the prediction horizon extends, thereby limiting the model’s capability.
In this work, we propose a novel model named ISMRNN to address the issues associated with SegRNN. Specifically, ISMRNN introduces an implicit segmentation structure that decomposes and maps the time series into encoded vectors through two linear transformations. This method facilitates more continuous processing during segmentation and enhances the utilization of information between different segments. Additionally, ISMRNN incorporates a residual structure with a linear layer, allowing some information to bypass the recurrent encoding structure, thus reducing information loss within the recurrent framework. Furthermore, we employ the Mamba structure(Gu and Dao, 2023) for preprocessing the time series, which aids in capturing long-term dependencies more effectively. The main contributions of ISMRNN can be summarized as follows:
@@
Incorporating the Mamba structure to improve information preprocessing.
The residual structure reduces information loss within the recurrent structure.
@@ 2 Problem Statement
Long-term time series analysis and forecasting comprise two critical components: (1) Trend identification and analysis, which entail the recognition of long-term directional changes within the data. This component is crucial for understanding the behavior of time series, enabling the comprehension of past patterns and the prediction of potential future developments. (2) Uncertainty and noise management: Due to the extended time range of long-term time series data, they often exhibit increased levels of uncertainty and noise. Effectively managing and mitigating these uncertainties and noises is vital for improving the accuracy of predictions.
@@ 2.1 Transformer Models
The core of applying Transformer models to long-term time series tasks lies in their ability to capture long-term dependencies using the attention mechanism (Hao and Cao, 2020). The general form of the self-attention mechanism(Vaswani et al., 2017) is given by:
Significant research efforts have advanced the application of Transformer architectures in long-term time series analysis. For instance, PatchTST(Nie et al., 2023) introduces a method of segmenting time series into patches, reducing computational complexity and improving the model’s ability to capture local semantic information. It also employs self-supervised learning, masking random patches and training the model to reconstruct them, enhancing forecasting accuracy. iTransformer(Liu et al., 2024) addresses scalability by integrating hierarchical attention mechanisms and temporal convolutional networks, capturing both short-term and long-term dependencies more effectively. This lightweight architecture balances accuracy and efficiency, making it suitable for real-world applications. Crossformer (Zhang and Yan, 2022) employs a Dimension-segment-wise structure to embed time series into a two-dimensional vector, effectively capturing dependencies across time and dimensions through its Two-stage-attention structure. The segmentation technique utilized in Crossformer is particularly advantageous for long-term time series forecasting.
@@ 2.2 MLP Models
TiDE (Das et al., 2023) introduces a residual structure within the linear layer to encode past time series and covariates, subsequently decoding the encoded time series with future covariates to obtain prediction results. The core operation of TiDE involves projecting features and encoding them through residual blocks, which can be described as:
where and denote the weights and biases of the projection layer, respectively. The ResBlock operation incorporates residual connections to preserve the integrity of the encoded features while integrating the covariates.
MLP-based methods for long-term time series analysis have demonstrated outstanding predictive accuracy while maintaining low time-space complexity, thus significantly promoting progress in the field of long-term time series forecasting.
@@ 2.3 CNN Models
The core of time series analysis based on CNN architectures lies in utilizing convolutional kernels to extract temporal and channel features. MICN (Wang et al., 2022) aims to fully exploit the latent information in time series by designing a multi-scale hybrid decomposition module to decompose the complex patterns of input time series. Each module within this design leverages subsampling convolution and equidistant convolution to respectively extract local characteristics and global correlations. TimesNet (Wu et al., 2023) transforms one-dimensional time series into a set of two-dimensional tensors based on multiple periodicities. It further introduces the TimesBlock to adaptively uncover these multi-periodic patterns and effectively extracts intricate temporal dynamics from the transformed tensors through parameter-efficient initializations. These methodologies have attained remarkable results in the domain of long-term time series forecasting.
2.4 RNN Models
The RNN structure is specifically designed for sequence tasks due to its ability to manage temporal dependencies through the recursive updating of hidden states. This mechanism enables the RNN to retain prior information and dynamically adjust internal states to capture short-term dependencies effectively. However, due to gradient issues, RNNs face challenges in capturing long-term dependencies. As the only RNN-based model specifically designed for long-term time series tasks, SegRNN has demonstrated exceptional performance. SegRNN addresses the challenge of retaining sequence information within the recurrent structure while minimizing information loss through two key methods:
SegRNN replaces point-wise iterations with segment-wise iterations. This approach balances the trade-off between retaining sequence information and the number of recurrent iterations, leading to improved performance.
SegRNN substitutes recurrent multi-step forecasting with parallel multi-step forecasting. This parallel decoding approach enhances the model’s predictive capability and inference speed.
By combining these two strategies, SegRNN achieves outstanding performance within a lightweight model structure and exhibits superior inference speed. However, the fixed segmentation approach of SegRNN limits the exchange of information between different segments and disrupts data continuity. Additionally, segmentation methods still face the challenge of information loss within the recurrent structure.
@@ 3.2 Implicit Segmentation
The segmentation method employed by SegRNN is implemented by truncating the time dimension. For a given time series , selecting a segmentation length results in segmented vectors of size , where represents the number of segments.
@@
To address these issues, we propose an implicit segmentation strategy, as illustrated in Figure 2. This strategy initially decomposes the time series into through linear mapping and subsequently transforms the information into the embedded hidden state . Notably, after segmentation, each segment includes both the information of that segment and some additional information, ensuring more continuous information processing and thereby enhancing the model’s performance.
@@ 3.3 Residual Structure
Extensive experiments have demonstrated that increasing the number of iterations within the recurrent structure can negatively impact the model’s information retention. The use of segmentation techniques can reduce the number of iterations in the recurrent structure from to . Specifically, for the vector obtained through segmentation in SegRNN, , we transform it in the RNN structure with a hidden layer dimension using the following formula:
@@
However, relying solely on segmentation does not fundamentally address the issue of information loss within the recurrent structure, which may become more pronounced as the prediction length increases. To address this issue, we incorporate a residual structure into the model. For the vector obtained through the decomposition by the first linear layer , we apply a linear mapping to , resulting in . Finally, we add to the corresponding position of the RNN encoder’s output vector to obtain the final encoder output vector .
@@
3.4 Mamba structure
The Mamba architecture excels in various fields, including large language models and long-term time series forecasting, due to its selective state space models (SSMs) that enable content-based parameter adjustment. This feature allows the architecture to dynamically select and process relevant information, making it highly effective for distinguishing signals from noise, particularly in time series forecasting.
We incorporate the Mamba architecture at the beginning of our workflow to preprocess the time series data. Specifically, we use the SSM component, which adjusts parameters based on the input signal. This setup enhances the model’s ability to selectively propagate and forget information along the sequence, focusing on relevant data points to improve forecasting accuracy.
Notably, we deviate from the standard Mamba architecture by excluding the convolutional layer, except for weather datasets. This choice is based on performance considerations, as we believe segmentation techniques effectively capture local information without the added computational load of convolutional processing.
@@ 4 Experiments
Our experiments are structured as follows. We first introduce the datasets used for long-term time series analysis. Subsequently, we provide a detailed description of the experimental setup and baseline models for comparison. Then we presents and analyzes the performance metrics. We further conducted ablation studies to investigate the effectiveness of the model. Finally, we evaluated the model’s efficiency to demonstrate its high spatiotemporal efficiency. All experiments in this section were conducted on two NVIDIA T4 GPU.
@@ 4.1 Datasets
The evaluation is conducted across six real-world datasets spanning various domains, including power transformer temperature, weather, electricity load, with channel ranging from 7 to 321 and frequencies from every 10 minutes to hourly recordings. Detailed dataset information is provided in Table 1.
@@ 4.3 Main Result
We select two evaluation metrics, MSE and MAE, with prediction steps of 96, 192, 336, and 720 time steps. The resulting forecasts are presented in Table 2. Note that all data, except for our method, originate from official code repositories and original papers.
@@ 4.4 Ablation study
We further conduct an ablation study to verify the effectiveness of incorporating the Mamba structure, implicit segmentation, and residual structures into our model. We performed experiments excluding the Mamba structure and implicit segmentation structure respectivly. For clarity, in the table below, "M" denotes preprocessing using the Mamba structure, and "LR" represents the addition of implicit segmentation and residual structure. The results are shown in Table 3.
@@
The ablation study reveals that the model incorporating the Mamba structure, along with implicit segmentation and residual structures performs the best overall. This is followed by the model that includes only implicit segmentation and residual structures. Interestingly, merely adding the Mamba structure does not significantly enhance model performance. This may be due to a synergistic effect between the two structures. Consequently, the ablation study confirms the superiority of our proposed structure.
@@ 4.5 Model efficiency
Compared to the SegRNN approach, our method introduces a slight increase in memory usage and training time, primarily due to the additional linear layers and the integration of the Mamba structure during the encoding phase. However, when compared to mainstream transformer-based models such as iTransformer and PatchTST, our model still demonstrates significant spatiotemporal advantages.
@@ 5 Conclusion
In this study, we introduce the ISMRNN model, an implicitly segmented RNN method that integrates the Mamba structure, implicit segmentation, and residual structure. The Mamba structure aids the model in capturing long-term dependencies, addressing the gradient vanishing and long-term dependency issues inherent in RNNs. Implicit segmentation resolves the fixed segmentation problem of SegRNN by maintaining sequence continuity, while the residual structure in the model reduces information loss associated with the recurrent structure. These enhancements significantly improve performance in long-term time series forecasting while keeping memory usage and training time relatively low. Experiments on six real-world benchmark datasets validate our findings. Notably, the proposed implicit segmentation method is adaptable to various scenarios beyond the specific long-term time series forecasting problem. Future work will explore how these structures enhance model performance and identify the specific attributes of time series data where the ISMRNN method excels.